### v1

# Beyond the Surface: Enhancing LLM-as-a-Judge Alignment with Human via Internal Representations

Abstract
The growing scale of evaluation tasks has led to the widespread adoption of automated evaluation using large language models, a paradigm known as "LLM-as-a-judge." However, improving its alignment with human preferences without complex prompts or fine-tuning remains challenging. In this work, motivated by preliminary findings that middle-to-upper layers encode semantically and task-relevant representations that are often more aligned with human judgments than the final layer, we propose lager, a lightweight and efficient framework for enhancing LLM-as-a-Judge alignment with human scoring, via internal representations. lager produces fine-grained judgment scores by aggregating cross-layer score-token logits and computing the expected score from a softmax-based distribution, with the LLM backbone kept frozen. lager fully leverages the complementary information across different layers, overcoming the limitations of relying solely on the final layer. We evaluate our method on the standard alignment benchmarks Flask, HelpSteer, and BIGGen using Spearman correlation, and find that lager achieves improvements of up to 7.5% over the best baseline across these benchmarks. Without reasoning steps, lager matches or outperforms reasoning-based methods. Experiments on downstream applications, such as data selection and emotional understanding, further show the effectiveness of our method.
1 Introduction
Large language models (LLMs) have witnessed significant progress in various tasks like math reasoning (Ying et al., 2024; Duan & Wang, 2024) and open-domain question answering (Allemang & Sequeda, 2024),and exhibit great generalization and reasoning abilities on unseen tasks and domains (Li et al., 2024c; Ye, 2024; Ni et al., 2024), which makes it possible to evaluate the model generations with an LLM (Bavaresco et al., 2024; Zheng et al., 2023). LLM-as-a-judge (Lin & Chen, 2023; Li et al., 2024a; Wu et al., 2024; Shiwen et al., 2024), an emerging approach to text evaluation, uses large language models to perform automated and scalable assessment of textual responses, reducing the reliance on human annotation. This paradigm has found widespread applications in various scenarios such as model evaluation (Lin & Chen, 2023), data synthesis (Kim et al., 2024a; Cui et al., 2023b; Wang et al., 2024b), and model enhancement via verification and critic agents (Pace et al., 2024; Badshah & Sajjad, 2024).
Researchers have explored various approaches to improve the consistency of judgment with human experts. Prompt-based methods (Lin & Chen, 2023; Liu et al., 2023) improve the judge performance with guidelines encouraging step-by-step reasoning (Chen et al., 2024; Lin & Chen, 2023; Zhuo, 2024) or carefully curated examples (Song et al., 2025). However, these methods rely on chain-of-thought reasoning, which introduces additional computational cost (Gu et al., 2024). Other researchers propose to finetune a LLM with a specialized dataset (Ye et al., 2024a; Cui et al., 2023a)to adapt the LLM to the judgment task. Although effective, these methods face the generalization problem (Doostmohammadi et al., 2024) as they have been adapted to the domain of the training data.
In this paper, we propose an effective LLM-as-a-judge framework called lager, to enhance LLM-as-a-Judge alignment with human via internal representations. Inspired by prior studies (Wang et al., 2020; Yang et al., 2022; Roh et al., 2024) showing that middle-to-upper layers encode richer semantic and task-specific information, we propose to aggregate the logits of score tokens from different layers. By applying a weighted combination followed by a softmax, we obtain a probability distribution over candidate scores, from which the final score is computed as the expected value. lager leverages the semantic diversity across different layers, allowing the final score to integrate both low-level lexical cues and high-level reasoning signals, thereby providing a more comprehensive reflection of the model’s understanding of the scoring task. Unlike the finetuning-based method, the parameters of the general LLM backbone are always kept frozen. lager outperforms baselines on alignment benchmarks such as Flask, HelpSteer, and BIGGen, with improvements of up to 7.5% in Spearman correlation. Although reasoning-based evaluation methods provide more transparency by exposing intermediate reasoning steps, their reasoning is often shallow and fails to improve reliability, leading to inferior performance. In contrast, without explicit reasoning, lager achieves comparable or superior results to these reasoning-based baselines. We further select instruction data with lager method and achieve better performance on the AlpacaEval-2.0 benchmark than various baselines. Experiments on more downstream applications show the effectiveness of the lager method.  [^1] Our code will be publicly available at https://github.com/sustech-nlp/LAGER.
Figure 1: (a): Vanilla scores may overlook meaningful distinctions by relying solely on the most probable score token. Better evaluation requires leveraging deeper signals beyond surface-level outputs. (b): The framework of lager. lager fully considers candidate score probabilities and aggregates cross-layer logits, enabling LLMs to act as robust evaluators with probabilistic scoring for reliable assessment.
2 Preliminary
A decoder-only LLM $f$ consists of an embedding layer $f_{\text{embd}}$, $L$ decoder layers $f_{\text{decoder}}$, and an output projection layer $f_{\text{unembd}}$. Given an input instruction $x$, the model generates a response $y$ through the following computation:
$y=f(x)=f_{\text{unembd}}\circ\underbrace{f_{\text{decoder}}^{(L)}\circ\cdots\circ f_{\text{decoder}}^{(1)}}_{\text{all decoder layers}}\circ f_{\text{embd}}(x),$ (1)
where $\circ$ denotes function composition, i.e., $(f\circ g)(x)=f(g(x))$. To predict the next token $x_{i}$, the model typically relies on the hidden state produced by the final decoder layer, denoted as
$\mathbf{h}^{(L)}_{i}=f_{\text{decoder}}^{(L)}\circ\cdots\circ f_{\text{decoder}}^{(1)}\circ f_{\text{embd}}(x_{<i}),$ (2)
where $x_{<i}=(x_{1},\dots,x_{i-1})$ represents the sequence of previous tokens. The output logits for token $x_{i}$ are then computed as:
$\hat{\mathbf{z}}_{i}=f_{\text{unembd}}(\mathbf{h}^{(L)}_{i})$. (3)
When incorporating a general LLM as the judge for a model response, the prompt template includes the evaluation task description $x_{d}$, user instruction $x_{i}$, the response to be evaluated $x_{r}$, and the scoring criteria and output format requirement $x_{c}$ (Liu et al., 2023, 2024; Hu et al., 2024a). We denote the full input as $X=\{x_{d},x_{i},x_{r},x_{c}\}$, and the set of candidate scores as $\mathbb{S}\subset\mathcal{V}$, where $\mathcal{V}$ is the model’s vocabulary. Let $\mathbf{h}^{(L)}_{n}$ denote the final-layer hidden state at the position where the score token is to be generated (e.g., after ‘Score:’). The model produces output logits:
$\hat{\mathbf{z}}_{n}=f_{\text{unembd}}(\mathbf{h}^{(L)}_{n})$,
which are then passed through a softmax over the model’s vocabulary $\mathcal{V}$:
$P_{L}(t|X,y_{<n})=\frac{\exp(\hat{\mathbf{z}}_{n}[t])}{\sum_{t^{{}^{\prime}}\in\mathcal{V}}\exp(\hat{\mathbf{z}}_{n}[t^{\prime}])}.$ (4)
The score token with the highest probability is the final score, referred to as the vanilla score:
$s_{L}^{*}=\arg\max_{s\in\mathbb{S}}P_{L}(s|X,y_{<n}).$ (5)
Figure 1(a) shows that vanilla score overlooks informative score distributions by relying solely on the top-probability score token, limiting its ability to distinguish response quality. More reliable and fine-grained evaluation requires the use of deeper internal representations beyond surface output.
3 Methods
3.1 Motivation
Figure 2: Agreement Between Human Ratings and Internal Layer Scores of Different Models.
Hidden representations across different layers have been widely observed to exhibit distinct characteristics: the bottom layers focus more on local lexical information, while the middle and top layers focus more on semantic and global information(Wang et al., 2020; Yang et al., 2022; Roh et al., 2024; Sun et al., 2025). Previous works (Sun et al., 2025; Alabi et al., 2024) using techniques such as PCA and shared-space analysis have shown that middle-layer hidden states are highly consistent and interchangeable across language adaptations, suggesting a shared representation space. Thus, they can be directly transformed to logits with the shared LLM output unembedding layer $f_{\text{unembd}}$.
Previous works on LLM-as-a-Judge predominantly rely on the final-layer hidden state, while ignoring the intermediate representations across layers. Hence, we investigate the judge scores derived from hidden representations at different layers to better understand the layer-wise evaluation behavior. Specifically, the logits at the $l_{th}(l<L)$ layer can be computed as follows:
$\hat{\mathbf{z}}_{l}=f_{\text{unembd}}(f_{\text{decoder}}^{(l)}\circ\cdots\circ f_{\text{decoder}}^{(1)}\circ f_{\text{embd}}(x_{<n}))=f_{\text{unembd}}(\mathbf{h}^{(l)}_{n})$, (6)
Following Equation 4, we can obtain the probability distribution $P_{l}(t|X,y_{<n})$over the vocabulary for each token, which allows us to compute the final judge score at the $l_{th}$ layer:
$s^{*}_{l}=\arg\max_{s\in\mathbb{S}}P_{l}(s|X,y_{<n}).$ (7)
Following the above setup, we conduct evaluations of different models’ layer-wise scoring performance across multiple benchmarks and analyzed their alignment with human judgments. The results [^2] For certain layers, if the calculated scores are identical across all samples, the Spearman correlation cannot be computed; such cases are shown as missing points in the figure. in Figure 2 show a consistent pattern across various benchmarks and models: one or more intermediate layers yield better agreement with human evaluations than the final layer.
In this paper, we focus on point-wise LLM evaluators without reference response. Inspired by our preliminary observations, we explore incorporating the hidden representations from intermediate layers to improve the judge score estimation. We propose the lager framework to enhance the consistency with judgment from human experts while maintaining the generalization ability of the LLM evaluator, and achieve better alignment with human score distributions.
3.2 Score Estimation Based on LLM Hidden Representations
For improved formality and clarity, we rewrite the LLM output unembedding layer $f_{\text{unembd}}$ as $\bm{W}_{\mathrm{unembd}}$, so $\hat{\bm{z}_{l}}$ can rewrite as:
$\hat{\bm{z}_{l}}=(f_{\text{decoder}}^{(l)}\circ\cdots\circ f_{\text{decoder}}^{(1)}\circ f_{\text{embd}}(x_{<n}))\bm{W}_{\mathrm{unembd}}=\mathbf{h}^{(l)}_{n}\bm{W}_{\mathrm{unembd}}.$ (8)
Based on the results in Figure 2, identifying consistently well-performing layers is challenging, as the layers exhibit varying levels of alignment with human judgments across different datasets. To address this, we first aggregate the logits corresponding to the set of candidate scores $\mathbb{S}$ from different layers using a set of weights $\bm{w}=\{w_{0},w_{1},\cdots,w_{L}\}$, where $w_{0}$ corresponds to the output of the embedding layer. The above procedure can be represented as:
$\hat{\bm{z}}=\sum_{i=0}^{L}w_{i}\hat{\bm{z}_{i}}=\sum_{i=0}^{L}w_{i}\mathbf{h}^{(i)}_{n}\bm{W}_{\mathrm{unembd}},$ (9)
following Equations 4 and 5, we can calculate a integer judge score. However, as illustrated in Figure 1(b), instead of relying on a single token, the full probability distribution over score candidates can be leveraged to capture richer evaluative information. This yields a probability distribution over the discrete score set, from which we compute a continuous, fine-grained score by taking the expected value. Therefore, we first extract the logits corresponding to all candidate score tokens from the full vocabulary logits and then aggregate them:
$\hat{\bm{z}}_{[\mathcal{M}]}=\sum_{i=0}^{L}w_{i}\hat{\bm{z}_{i}}_{[\mathcal{M}]}=\sum_{i=0}^{L}w_{i}[\mathbf{h}^{(i)}_{n}\bm{W}_{\mathrm{unembd}}]_{\mathcal{M}},$ (10)
where $\mathcal{M}=\{\mathrm{Tokenize}(s)|s\in\mathbb{S}\}$ denotes the set of tokenized sequences from $\mathbb{S}$. The final probability distribution of $\mathbb{S}$ is then obtained by applying a softmax over the aggregated logits:
$P(s)=\frac{\exp(\hat{\mathbf{z}}[s])}{\sum_{s^{{}^{\prime}}\in\mathbb{S}}\exp(\hat{\mathbf{z}}[s^{\prime}])},\quad s\in\mathbb{S}$ (11)
by taking the expectation over the score distribution, we can obtain the final fine-grained judge score:
$s^{*}=\mathbb{E}_{s\sim P(s)}[s]=\sum_{s\in\mathbb{S}}s\times P(s)$ (12)
Normalizing before aggregation removes the relative scale information of logits and overemphasizes less important layers. We compare both settings in the ablation study (Section 4.3).
3.3 Lightweight Training of Layer Weights
We propose two types of layer weights $\bm{w}$. One is to apply average aggregation $w_{l}=1/(L+1)$ (denoted as lager (w.o. tuning) in Table 1), the other is to tune the lightweight $L+1$ parameters with a small-scale validation set (denoted as lager (w. tuning)). With a frozen backbone and minimal learnable parameters, we refer to the dataset as a validation set to distinguish it from finetuning-based LLM evaluators. To enhance the model’s performance while aligning its predictions more closely with the distribution of human scores, we adopt a combination of cross-entropy(CE) loss and mean absolute error (MAE) loss, balanced by a weighting hyperparameter $\alpha$.
$\mathcal{L}_{\text{Final}}$ $=\alpha\cdot\mathcal{L}_{\text{CE}}+(1-\alpha)\cdot\mathcal{L}_{\text{MAE}}$ (13)
$=\alpha\cdot\left(-\frac{1}{\mathcal{|B|}}\sum_{i=1}^{\mathcal{|B|}}\sum_{s=1}^{\mathbb{S}}\mathbb{I}(s=s_{\mathrm{truth}}^{i})\log P_{i}(s)\right)+\left(1-\alpha\right)\cdot\left(\frac{1}{2\mathcal{|B|}}\sum_{i=1}^{\mathcal{|B|}}(s_{i}^{*}-s_{\mathrm{truth}}^{i})^{2}\right)$
where $\mathcal{B}$ is the batch of samples in a training step, and $s_{truth}^{i}$ is the human-annotated score for each sample. Please refer to Appendix E.5 for specific training settings and details. It is important to note that the weights are tuned only once for each backbone and subsequently applied across all benchmarks. As the LLM remains frozen, no additional judge model is needed for scenarios like self-improvement. In contrast, finetuning-based methods require domain-specific judge models and suffer from limited generalization due to scarce labeled data.
3.4 Discussion: Applicability under Restricted Access Settings
While our proposed framework relies on accessing the intermediate layers of the LLM, it can naturally adapt to more restricted scenarios. For instance, when only final-layer logits are available (e.g., API-based models), we can extract the logits corresponding to score tokens, apply softmax, and compute the expected score. The method proposed in prior work (Liu et al., 2023) generates a continuous score estimate by avoiding argmax on discrete tokens. We refer to this setup as expectation score (E-Score). The evaluation uses the vanilla score, based on the most likely token, when only the final token prediction is observable. Both settings can be viewed as simplified special cases within our framework. We treat both scores as comparable baselines in our experiments.
4 Experiments
4.1 Experiments Setup
Benchmarks
Due to potential preference leakage (Li et al., 2025) between models, we evaluate our method on human-annotated datasets for reliable results. We chose three point-wise benchmarks to ensure a comprehensive evaluation: Flask (Ye et al., 2024a), HelpSteer (Wang et al., 2023), and BiGGen Bench [^3] https://huggingface.co/datasets/prometheus-eval/BiGGen-Bench-Results  (Kim et al., 2024b). In these benchmarks, human-annotated scores range from 1 to 5. Please refer to Appendix E.2 for the details.
Models
We experiment with multiple backbone models of varying sizes and from different families, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), InternLM3-8B-Instruct (Cai et al., 2024), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), Qwen-2.5-14B-Instruct (Qwen et al., 2025), Mistral-Small-24B-Instruct and LLAMA-3.3-70B-Instruct.
Baselines
We compare lager with three baselines: GPTScore, vanilla score (VScore) and expectation score (E-Score). At the same time, we compare the commonly used API-based model GPT-4o-mini. We also experiment with methods that explicitly train the LLM backbones, namely TIGERScore-7B and Prometheus2-7B. While lager does not modify the backbone models, these methods are included solely for reference and are not directly compared. Please refer to Appendix E.3 for specific details about the baselines.
Evaluation Metrics
Following previous works (Gu et al., 2024; Bai et al., 2023; Liu et al., 2025), we use the commonly adopted Pearson and Spearman correlations to measure the consistency between LLM-predicted scores and human annotations. See Appendix E.4 for further details.
Evaluation details
We adopt greedy decoding for its stability and consistency in evaluation (Lin & Chen, 2023). To comprehensively assess the effectiveness of our method, we conduct evaluations under two different conditions: 1) Direct: Directly providing a score during evaluation without reasoning. 2) Reasoning: Previous work (Liu et al., 2023; Wang et al., 2024a; Hu et al., 2024a) shows that reasoning improves judgment quality, with reasoning-first setups performing best(Chen et al., 2024; Murugadoss et al., 2024; Gu et al., 2024). Thus, we adopt a reasoning-first approach in this study. Please refer to Appendix G for the prompt templates. Unless otherwise specified, we highlight the best results in each table using bold, and denote the second-best with underlining.
Table 1: Spearman Correlation Results (Part 1 of 2): lager vs. Baselines on Flask, HelpSteer, and BiGGen Bench. See Table 5 for the remaining results. These results are statistically significant, with p-values less than 1e-5. See Table 6 for the full Pearson correlation results.
Model Flask HelpSteer BIGGen Bench Average
Direct Reasoning Direct Reasoning Direct Reasoning
Fine-tuned Models
TIGERscore-7B - 0.175 - 0.118 - 0.171 0.155
Prometheus2-7B - 0.413 - 0.514 - 0.367 0.431
Close-source Model via API
GPT-4o-mini
Vscore 0.526 0.535 0.482 0.535 0.534 0.509 0.520
E-Score 0.579 0.561 0.500 0.541 0.573 0.530 0.547
Open-source Models
InternLM3-8B-Instruct
GPTScore -0.087 - -0.062 - -0.257 - -0.135
Vscore 0.423 0.449 0.388 0.425 0.374 0.441 0.417
E-Score 0.515 0.472 0.453 0.430 0.470 0.470 0.468
lager (w.o tuning) 0.449 0.468 0.426 0.429 0.374 0.469 0.436
lager (w. tuning) 0.545 0.489 0.515 0.474 0.507 0.49 0.501
Qwen-2.5-14B-Instruct
GPTScore 0.001 - -0.014 - -0.142 - -0.052
Vscore 0.547 0.537 0.420 0.423 0.458 0.461 0.474
E-Score 0.579 0.555 0.447 0.452 0.502 0.457 0.499
lager (w.o tuning) 0.572 0.567 0.433 0.473 0.503 0.507 0.509
lager (w. tuning) 0.612 0.572 0.443 0.472 0.567 0.524 0.531
Mistral-Small-24B-Instruct
GPTScore 0.016 - -0.008 - -0.147 - -0.046
Vscore 0.528 0.505 0.420 0.459 0.542 0.533 0.498
E-Score 0.577 0.532 0.442 0.486 0.585 0.555 0.530
lager (w.o tuning) 0.591 0.542 0.452 0.485 0.589 0.562 0.537
lager (w. tuning) 0.596 0.542 0.449 0.487 0.598 0.566 0.540
4.2 Main Results
Tables 1 and 5 report Spearman results of lager compared to baselines across three benchmarks. Section 5.3 further analyzes the performance variation of lager and baseline methods across different model scales, under the setting where models from the Qwen2.5 family are used as frozen backbones, and compares their results on the same set of benchmarks.
The Effectiveness and Scalability of LAGER
lager with tuning consistently outperforms all non-training baselines across diverse LLM backbones on all three benchmarks. This highlights the importance of leveraging internal logits and output distributions for accurate scoring. Although Mistral-Small-24B-Instruct and Qwen2.5-14B-Instruct underperform GPT-4o-mini on Vscore, they achieve comparable results when using lager. GPTScore performs poorly due to its reliance on generation probabilities, which overlook the semantic and logical quality. Due to its focus on error detection, the fine-tuned TIGERscore-7B shows relatively limited performance. Compared to the fine-tuned Prometheus2-7B, lager also enables InternLM3-8B-Instruct and LLaMA3.1-8B-Instruct(results shown in Table 5), two models of comparable size, to surpass it even though they initially have weaker performance on Vscore.
The Robustness of LAGER with Tuning
The average E-Score performance of models like Qwen-2.5-14B-Instruct, LLAMA-3.3-70B-Instruct, and Mistral-7B-Instruct-v0.3 is sometimes lower than VScore. Specifically, the average E-score performance of LLAMA-3.3-70B-instruct (results shown in Table 5) across all benchmarks is 2.4 points lower than Vscore. In contrast, lager with tuning achieves almost the best performance across all settings and benchmarks, consistently outperforms Vscore. lager can closely match human score distribution and incorporate more fine-grained information. More details are discussed in Section 5.
Reasoning Is Not Always Better
The richness of information aggregation increases gradually from Vscore, E-Score to lager (with tuning). Figure 6 shows that reasoning leads to an improvement in Vscore performance, a divergence in E-Score performance (with half of the models improving and the other half declining), and consistently weaker results for lager (with tuning) under the reasoning setting compared to direct. We speculate that this phenomenon may be due to reasoning potentially causing the model to become overconfident (Yang et al., 2024; Wagner et al., 2024), which in turn leads to inaccurate evaluation scores.
4.3 Ablation Study
Table 2: Ablation Study Results on FLASK and HelpSteer Datasets Evaluated by Spearman Corr.
Strategies For lager InternalLM Mistral
ID Exp. Max. Logits agg. Prob. agg. Tuning Flask HelpSteer Flask HelpSteer
$\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ 0.545 0.515 0.347 0.403
$\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 0.452 0.448 0.299 0.344
$\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 0.373 0.374 0.268 0.264
$\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 0.367 0.379 0.268 0.267
$\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ 0.449 0.426 0.338 0.401
$\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ 0.442 0.445 0.306 0.358
$\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ 0.379 0.367 0.265 0.264
$\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 0.364 0.363 0.268 0.267
$\checkmark$ $\times$ $\times$ $\times$ $\times$ 0.515 0.453 0.239 0.296
[…]
To better understand each design choice of lager, we conduct a comprehensive ablation study on the direct setup using Mistral-7B-Instruct-v0.3 and InternLM3-8B-Instruct models. Table 2 shows the results of the Spearman correlation. The expectation (Exp.) score and the maximum (Max.) score are two alternative designs. Given the score distribution of a certain layer, we use Eqn.12 to calculate the expected score or use the score with the highest probability.
Logits agg. and prob. agg. are two alternative designs, aggregating information from all hidden layers at the logits level or output distribution level. Removing logits agg. and prob. agg. indicates that we do not use internal states. It is observed that the full configuration (Exp. score + Logits agg. + Tuning) achieves the best performance across all settings, with a maximum Spearman correlation of 0.545; Fine-tuning brings up to +0.10 improvement; expectation scoring outperforms maximum scoring (up to +0.17); and logits aggregation surpasses probability aggregation (up to +0.07). Multi-layer integration is generally effective, especially for Mistral; however, in untuned settings, InternalLM occasionally performs better without aggregation, indicating model-specific differences.
Figure 3: Analysis of Llama3.1-8B-Instruct and Mistral-7B-Instruct-V0.3 using HelpSteer and Flask. (a) and (b) illustrate the performance of E-scores computed from different layers. (c) depicts the average cosine similarity heatmap of hidden states across layers.
5 Analyses
5.1 Understanding Internal States
Having observed that internal states contribute significantly to the success of lager, we now aim to understand how internal states in each layer contribute to the judgment score. Specifically, we apply lager to each layer of the backbone LLM, namely using the expected score from the output distribution of each layer, and report the consistency with human scores. Figure 3 illustrates the results of the Mistral-7B-Instruct-v0.3 and LLaMA-3.1-8B-Instruct backbones. We observe the following: (1)The bottom layers exhibit low or even negative correlation with human scores, indicating their limited utility in evaluation tasks. (2)The middle-upper layers (approximately layers 20 to 30) show the highest correlation with human scores, suggesting they carry the most informative representations for judgment. However, there is a drop in correlation at the topmost layer, implying that certain judgment-relevant signals may be diluted or lost in the final transformation stages. Thus, relying solely on the top layer is insufficient for precise evaluation.
To better understand the differences between internal and top layers, we visualize the cosine similarity between hidden states across layers at the position of the score token using the Flask and HelpSteer datasets. Based on Figure 3(c), the similarity between layers generally decreases as the distance between layers increases. For the bottom layers, the representation changes rapidly with increasing layer number. In contrast, the representation changes relatively slowly for the middle-upper layers, and the representations across these layers remain similar over a broader range. The topmost layer exhibits low similarity to its neighboring layers, displaying a distinct pattern.
Based on the above observations, we believe that: The middle-upper layers exhibit relatively consistent representations and stronger alignment with human scores, indicating that these layers may deserve more attention. Meanwhile, the topmost layer shows comparatively weaker judgment ability, likely because its focus on next-token prediction leads to the loss of fine-grained evaluative information. Instead of searching for optimal layers, which vary across LLM backbones and are hard to identify, lager directly learns a weighted combination of all layers. This approach is both lightweight and effective, requiring only a small validation set and transferring well across benchmarks (Section 4.2, Appendix D.1, Appendix D.2) and tasks (Section 6.1, Appendix D).
5.2 LAGER Yields More Human-Aligned and Fine-Grained Score Distributions
Figure 4: The KDE Plot of Different Score Distributions, evaluated on Flask with LLaMA-3.1-8B-Instruct backbone.
To better understand how lager improves over both the VScore and the E-Score, we evaluate LLaMA-3.1-8B-Instruct on the Flask dataset to obtain the Vanilla Score, E-Score, and lager, and visualize the min-max normalized scores and human annotations using Gaussian kernel density estimation (KDE). Details of the KDE visualization procedure can be found in Appendix E.1. Figure 4 visualizes the score distributions. Compared to E-Score and VScore, lager yields a distribution that more closely matches human judgments, with noticeably higher overlap. lager mitigates the high-score bias of VScore and E-Score by shifting the score distribution to the left and producing more evenly distributed scores.
Table 3: Distribution distance of VScore and lager to human score, evaluated on Flask with LLaMA-3.1-8B-Instruct backbone.
LLaMA-3.1-8B-Instruct $\mathrm{D_{KL}}(\downarrow)$ $\mathrm{MSE}(\downarrow)$
VScore 0.312 0.112
E-Score 0.102 0.092
lager (w. tuning) 0.087 0.060
We also quantitatively compare lager, E-Score and VScore by evaluating the Kullback-Leibler (KL) divergence and the mean squared error (MSE). The KL divergence of a candidate score $S$ is computed as $\mathrm{D}_{\mathrm{KL}}(\tilde{p}_{hs}||\tilde{p}_{s})$ where $\tilde{p}_{hs}$ and $\tilde{p}_{s}$ represent the human score distribution and the candidate score distribution, respectively. The results, presented in Table 3, demonstrate that lager more accurately approximates the distribution of human-annotated scores.
5.3 From Tiny to Huge: LAGER Brings Improvements Across Model Scales
To further investigate whether lager can enhance human alignment across a broader range of model scales, we conducted experiments using three benchmarks on instruct models of varying sizes from the Qwen2.5 family, including 0.5B, 1.5B, 3B, 7B, 14B, 32B and 72B.
The Figure 5 and Figure 7 show the alignment effectiveness of Vanilla Score, E-Score, and lager (w. tuning) across Qwen2.5 models from 0.5B to 72B. Both direct and reasoning conditions show a mostly upward trend in Spearman correlation coefficients as model size increases across Flask, Helpsteer, and BIGGen datasets, with lager leading (e.g., 0.658 in Flask for direct, 0.598 for reasoning). When the model size exceeds 14B, lager achieves greater performance gains over the backbone model than the VScore computed using GPT-4o-mini. Notably, on the BIGGen dataset, lager shows a significant improvement on the 0.5B model scale, elevating the Spearman correlation from a negative value (-0.4 for the Vanilla Score) to a positive 0.1, marking a substantial gain compared to the baseline. These results demonstrate its effectiveness across different model scales and datasets.
Figure 5: Comparison of Spearman Results for Qwen2.5 Models Across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) (Direct Condition).
6 Applications of LAGER
6.1 Instruction Data Selection
Instruction data selection (Cao et al., 2024; Shen, 2024) involves choosing a subset of high-quality data from a given instruction dataset to enhance both the efficiency and performance of the tuning process. In this experiment, we use lager as a scoring metric to select a 10% subset from the alpaca-cleaned-52k dataset [^4] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023). We compare our subset, Highest lager, with four baselines on finetuning the LLaMA3-8B-base model: Longest instructions, Longest responses (Shen, 2024), Highest VScore, and SuperFiltering (Li et al., 2024d). SuperFiltering is an explicitly designed instruction data filtering method. For methods that require a backbone LLM, we utilize LLaMA3.1-8B-Instruct for a fair comparison. More details are in Appendix F.
Table 4: The Performance of LLaMA3-8B-base Fine-tuning with Instruction Data Subsets on AlpacaEval 2.0: Length-Controlled (LC) win rate against GPT-4-1106-Preview (805 Test Examples), evaluated with GPT-4o-mini.
Filtering criteria Ratio AlpacaEval-2.0
LC win rate average length
No filtering 100% 7.73 1070
Longest instructions 10% 6.92 946
Highest VScore 10% 9.42 1206
Highest E-Score 10% 11.50 1282
SuperFiltering 10% 10.69 1367
Longest responses 10% 11.82 1506
Highest lager (w.o. tuning) 10% 11.77 1325
Highest lager (w. tuning) 10% 12.65 1248
As shown in Table 4, our method with tuning achieves the best performance, outperforming the second-best method, the Longest responses, and the explicitly designed method, SuperFiltering, by 0.83 and 1.96 LC win rate, respectively. Even without tuning, our method outperforms most baselines, improving the VScore by a 2.35 LC win rate. And the highest lager (w tuning) has relatively shorter average lengths than most baselines. This indicates that our tuning model does not achieve better scores by generating longer responses, but by learning to follow instructions more effectively.
In addition to the application mentioned above, lager can be applied to other scoring-based tasks, such as LLM emotional understanding and LLM Recognizing its Knowledge Boundaries. lager shows strong transferability, even when there is a mismatch between the task’s scoring range and that of the data used for training. Please refer to Appendix D.1 and Appendix D.2 for detailed discussions.
7 Conclusion
In this work, we propose lager, an effective LLM-as-a-judge framework to estimate the judge score from the logits information from different LLM layers. lager offers training-free and tunable options, allowing flexible integration with various LLMs to deliver fine-grained judgment services. Unlike previous methods that explicitly update LLM backbones, lager provides a lightweight tuning solution that only trains weights on a small validation dataset, and the learned weights can be transferred across tasks. Experiments on three comprehensive alignment benchmarks demonstrate the effectiveness of lager. Specifically, lager outperforms all baselines that do not explicitly train LLM backbones. Additionally, lager demonstrates strong performance in improving alignment with human annotations across models of varying scales and families. Even without explicit reasoning, lager outperforms various reasoning-based baselines, including those specifically trained for reasoning tasks, while requiring fewer generation tokens and offering higher efficiency. lager is quite general and can be applied to various use cases, including but not limited to instruction data selection, recognizing its knowledge boundaries and emotional understanding.
Appendix A Impact Statement
Limitation
Our method relies on obtaining the hidden states from all model layers during the training and evaluation phases. This requirement means full internal access to the model’s architecture and intermediate representations is essential for optimization. Consequently, the method is applicable and can be optimized only for open-source models, where such internal states are accessible. In contrast, for closed-source or proprietary models, where access to internal hidden states is typically restricted, our method cannot be deployed to improve performance, limiting its applicability in those scenarios.
Ethics Statement
We affirm that this research adheres to the ethical standards in AI and machine learning. No personal, sensitive, or proprietary data was used in the experiments. All datasets employed are publicly available and widely used in the research community. Our code is publicly available, and all experiments are reproducible. We have carefully considered the potential societal impacts and are committed to applying the method responsibly, ensuring it has a positive and controllable impact on society.
Future Societal Consequences
Our approach relies on comprehensive access to the internal states of models, which has profound implications for the future ecological landscape of AI development. As open-source models continue to evolve, this method helps drive the democratization of performance optimization, enabling researchers to improve models transparently and collaboratively, fostering innovation, and enhancing the accountability of AI systems. However, it cannot be applied to closed-source or proprietary models, highlighting the growing gap between open and closed AI ecosystems and potentially increasing dependence on a few dominant providers. Furthermore, since this approach is primarily used for improving judgment models, there is a potential risk of misuse, such as interfering with or manipulating evaluation results, reminding us to be mindful of possible negative impacts when promoting its application, ensuring responsible and controlled use.
Appendix B Related Work
B.1 Evaluation of Text Generation Tasks
Over the past decades, automatic evaluation metrics, including statistical methods and embedding-based approaches, have played a crucial role in assessing the quality of generated text. The metrics BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) have long been standard for evaluating generated text. BERTScore (Zhang et al., 2020) evaluates the similarity between two sentences by using a pre-trained BERT model to generate contextual embeddings for each token and calculating the sum of their cosine similarities. BARTscore (Yuan et al., 2021) evaluates the quality of generated text by calculating the weighted logarithmic probability of one text given another text. Chung et al. (2023) found that these metrics cannot be accurately evaluated for semantically equivalent expressions but syntactically different. Therefore, trying to evaluate harder-generated text requires more efficient methods. GPTScore (Fu et al., 2023) evaluates text quality through conditional generation probabilities, enabling customizable assessments using natural language instructions. However, these methods all require reference answers for a stable evaluation.
B.2 LLM-as-a-Judge
The advancement of LLMs has made LLM-as-a-judge possible, where an LLM-based evaluator judges the model responses based on user instructions and the criteria from different aspects of the prompt. The evaluation is flexible as the evaluation criteria can be adjusted dynamically in the prompt instead of retraining the evaluator from scratch. The judgment can be interpretable and detailed criticism as well as feedback are generated before the judge score. The LLM-based evaluator is more efficient and scalable than human experts at much less expense. It is more effective and applicable than statistical evaluation metrics like BLEU and embedding-based metrics like BERTScore.
Now there are three different categories of LLM judges: The Point-wise judge model (Kim et al., 2024a; Lin & Chen, 2023; Dong et al., 2024) evaluates individual responses independently, scoring them based on predefined criteria; The pair-wise judge model(Li et al., 2024b; Kim et al., 2024b; Jung et al., 2024) compares two responses directly, determining which one is superior based on a set of evaluation dimensions; The List-wise judge model(Liu et al., 2025) ranks multiple responses in terms of quality, assigning a rank to each based on the overall evaluation. Researchers explore effective approaches to apply LLMs as judges by prompt-based or finetuning-based methods in various scenarios. The prompt-based approaches aim to enhance judgment via step-by-step instructions as well as multi-turn optimizations. G-Eval (Liu et al., 2023) evaluates NLG output quality using Chain-of-Thought (CoT) and Form-Filling paradigms. Portia (Li et al., 2024e) proposes to evaluate the responses with step-by-step reasoning guidelines. Active-critic (Xu et al., 2024) first generates evaluation criteria from data and iteratively refines them. However, the performance of prompt-based LLM evaluators is unsatisfactory (Gu et al., 2024). The finetuning-based approaches further optimize the LLMs with the specialized dataset to adapt the LLMs to judgment tasks either with instruction tuning (Ye & Ng, 2024; Li et al., 2024b) or preference tuning (Ye et al., 2024b; Hu et al., 2024b). InstructScore (Xu et al., 2023) combines human guidance with implicit knowledge from models like GPT-4, finetuned on the LLaMA model, to provide both scores and human-readable diagnostic reports. TIGERScore (Jiang et al., 2023b) is finetuned on the LLaMA-2 model to generate detailed error analysis, helping users understand each identified error and its associated penalty score. Prometheus (Kim et al., 2024a) and Prometheus 2 (Kim et al., 2024b) are finetuned for fine-grained evaluation, achieving new state-of-the-art performance on judgment benchmarks. In this work, we focus on the single-LLM point-wise evaluator without a reference response. Our proposed lager method improves judgment performance by estimation with aggregated layer-wise logits, which is different from the idea of prompt-based and finetuning-based methods.
Appendix C Addition Experimental Results
In this section, Figure 6 shows the comparison of the performance of different scoring methods across models on direct and reasoning conditions. Figure 7 shows the comparison of spearman correlation coefficients for Qwen2.5 models across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) under the reasoning condition. We present the Spearman correlation coefficients of lager and other baselines for LLaMA3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Mistral-Small-24B-Instruct on Flask, HelpSteer, and BiGGen Bench, as shown in Table 5. In Table 6, we present the Pearson correlation coefficients for all models, including lager and other baselines, on the Flask, HelpSteer, and BiGGen Bench.
Figure 6: Average performance comparison of different scoring methods across models on direct and reasoning tasks over three benchmarks.
Figure 7: Comparison of Spearman Correlation Coefficients for Qwen2.5 Models Across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) (Reasoning Condition)
Table 5: Spearman Correlation Results (Part 2 of 2): Continuation of Table 1. These results are statistically significant, with p-values less than 1e-5. See Table 6 for the full Pearson correlation results.
Model
Flask
Direct
Flask
Reasoning
HelpSteer
direct
HelpSteer
Reasoning
BIGGen Bench
Direct
BIGGen Bench
Reasoning
Average
Fine-tuned Models
TIGERscore-7B - 0.175 - 0.118 - 0.171 0.155
Prometheus2-7B - 0.413 - 0.514 - 0.389 0.448
Close-source Model via API
GPT-4o-mini
Vscore 0.526 0.535 0.482 0.535 0.534 0.509 0.520
E-Score 0.579 0.561 0.500 0.541 0.573 0.530 0.547
Open-source Models
LLaMA3.1-8B-Instruct
GPTScore 0.061 - -0.022 - -0.162 - -0.041
Vscore 0.334 0.429 0.374 0.518 0.273 0.390 0.386
E-Score 0.386 0.446 0.464 0.525 0.352 0.403 0.429
lager (w.o tuning) 0.472 0.456 0.520 0.524 0.475 0.443 0.482
lager (w. tuning) 0.477 0.460 0.515 0.524 0.482 0.444 0.484
Mistral-7B-Instruct-v0.3
GPTScore 0.258 - 0.209 - 0.183 - 0.217
Vscore 0.266 0.269 0.267 0.364 0.138 0.280 0.264
E-Score 0.239 0.279 0.296 0.380 0.185 0.283 0.277
lager (w.o tuning) 0.338 0.295 0.401 0.377 0.353 0.329 0.349
lager (w. tuning) 0.347 0.298 0.403 0.376 0.357 0.333 0.352
LLAMA-3.3-70B-Instruct
GPTScore 0.042 - 0.014 - -0.193 - -0.046
Vscore 0.518 0.567 0.435 0.494 0.559 0.539 0.519
E-Score 0.464 0.540 0.444 0.488 0.530 0.506 0.495
lager (w.o tuning) 0.610 0.598 0.506 0.520 0.597 0.585 0.569
lager (w. tuning) 0.611 0.598 0.504 0.519 0.602 0.584 0.570
Table 6: Pearson correlation coefficients of lager and other baselines on Flask, HelpSteer, and BiGGen Benchmarks. These results are statistically significant, with p-values less than 1e-5. These results are statistically significant, with p-values less than 1e-5.
Model
Flask
Direct
Flask
Reasoning
HelpSteer
direct
HelpSteer
Reasoning
BIGGen Bench
Direct
BIGGen Bench
Reasoning
Average
Fine-tuned Models
TIGERscore-7B - 0.227 - 0.173 - 0.116 0.172
Prometheus2-7B - 0.453 - 0.502 -
Close-source Model via API
GPT-4o-mini
Vscore 0.563 0.581 0.484 0.529 0.576 0.538 0.545
E-Score 0.588 0.587 0.503 0.532 0.588 0.543 0.557
Open-ource Models
LLaMA3.1-8B-Instruct
GPTScore 0.109 - 0.028 - -0.109 - 0.009
Vscore 0.393 0.486 0.376 0.517 0.329 0.441 0.424
E-Score 0.386 0.494 0.427 0.525 0.380 0.452 0.444
lager (w.o tuning) 0.487 0.514 0.483 0.525 0.473 0.481 0.494
lager (w. tuning) 0.493 0.512 0.486 0.524 0.485 0.482 0.497
Mistral-7B-Instruct-v0.3
GPTScore 0.336 - 0.166 - 0.154 - 0.219
Vscore 0.292 0.305 0.302 0.355 0.194 0.321 0.295
E-Score 0.307 0.313 0.333 0.363 0.221 0.329 0.311
lager (w.o tuning) 0.380 0.343 0.412 0.380 0.357 0.376 0.375
lager (w. tuning) 0.378 0.337 0.413 0.377 0.361 0.375 0.374
InternLM3-8B-Instruct
GPTScore -0.104 - -0.047 - -0.128 - -0.093
Vscore 0.468 0.516 0.374 0.429 0.399 0.475 0.444
E-Score 0.544 0.532 0.446 0.441 0.477 0.489 0.488
lager (w.o tuning) 0.475 0.517 0.417 0.441 0.371 0.471 0.449
lager (w. tuning) 0.568 0.551 0.494 0.474 0.497 0.501 0.512
Qwen-2.5-14B-Instruct
GPTScore -0.026 - -0.065 - -0.108 - -0.066
Vscore 0.585 0.581 0.428 0.420 0.494 0.507 0.503
E-Score 0.612 0.594 0.448 0.436 0.517 0.515 0.520
lager (w.o tuning) 0.612 0.600 0.430 0.436 0.496 0.503 0.513
lager (w. tuning) 0.645 0.618 0.443 0.457 0.576 0.551 0.548
Mistral-Small-24B-Instruct
GPTScore -0.020 - -0.044 - -0.088 - -0.051
Vscore 0.573 0.545 0.418 0.440 0.589 0.575 0.523
E-Score 0.608 0.561 0.442 0.463 0.624 0.589 0.548
lager (w.o tuning) 0.603 0.577 0.448 0.477 0.628 0.603 0.556
lager (w. tuning) 0.608 0.569 0.448 0.477 0.634 0.604 0.557
LLAMA-3.3-70B-Instruct
GPTScore 0.022 - -0.033 - -0.121 - -0.044
Vscore 0.544 0.607 0.435 0.482 0.590 0.587 0.541
E-Score 0.554 0.610 0.444 0.485 0.597 0.588 0.546
lager (w.o tuning) 0.627 0.634 0.506 0.515 0.644 0.624 0.592
lager (w. tuning) 0.626 0.630 0.503 0.512 0.645 0.620 0.589
Appendix D Addition Applications of LAGER
D.1 Emotional Understanding
LLMs have demonstrated strong capabilities to comprehend and interpret complex emotions and their meanings in social contexts. Therefore, beyond merely assessing the quality of responses, LLM can also be used as an emotion judge. In the widely used emotion understanding benchmark EQ-Bench (Paech, 2024), EQ-Bench is a benchmark for language models designed to assess emotional intelligence. It uses a specific question format in which participants are required to read a conversation and then rate the intensity of one character’s emotional response. Each question is explanatory and aims to assess the ability to predict the intensity of four different emotions. The dataset originally contained 117 questions, and after processing, we obtain 430 instances, each corresponding to one emotion per conversation. The emotional intensity annotated by humans ranges from 1 to 9, with higher scores indicating stronger emotions. The LLM is asked to predict the intensity of the emotional states of characters in a dialogue with a score ranging from 1 to 9. In this part, we evaluate lager on EQ-Bench and report the consistency with human evaluations.
As shown in Table 7, lager significantly outperforms the widely used vanilla score across all LLM backbones and consistency metrics, with an average improvement of 8.42 points without tuning and 9 points with tuning. This demonstrates that lager is effective not only for evaluating the quality of responses but also in other LLM-as-a-Judge scenarios, such as emotional understanding. lager is tuned on the HelpSteer validation set, which has a different scoring domain and range compared to EQ-Bench (ranging from 1-5 versus 1-9). This further highlights the robustness of lager with tuning, as it successfully transfers from one domain and score range to other domains with different judging prompts and score ranges.
Table 7: The Spearman and Pearson correlation coefficient measures the consistency between reference sentiment intensity and the scores from LLaMA3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen-2.5-14B-Instruct, on the processed EQ-Bench.
Model Scoring Type Processed-EQ-Bench
Pearson Spearman
Mistral-7B-Instruct-v0.3 VScore 0.56 0.596
E-Score 0.574 0.618
lager (w.o. tuning) 0.593 0.632
lager (w. tuning) 0.598 0.635
LLaMA-3.1-8B-Instruct VScore 0.456 0.478
E-Score 0.608 0.652
lager (w.o. tuning) 0.627 0.653
lager (w. tuning) 0.634 0.657
Qwen-2.5-14B-Instruct VScore 0.706 0.707
E-Score 0.739 0.759
lager (w.o. tuning) 0.745 0.758
lager (w. tuning) 0.756 0.763
D.2 Enhancing LLM in Knowing When They Don’t Know
Table 8: Self-knowledge Comparison of lager and baselines. The evaluation metric is the F1-Score.
Model Classification Methods
SimCSE VScore E-Score
lager
(w.o tuning)
lager
(w. tuning)
Mistral-7B-Instruct-v0.3 0.427 0.605 0.62 0.62 0.624
LLaMA-3.1-8B-Instruct 0.490 0.553 0.574 0.59 0.570
Knowing when they don’t know, also known as self-knowledge (Yin et al., 2023), is crucial for enhancing the reliability and trustworthiness of LLMs. In this section, we explore whether lager can improve an LLM’s self-knowledge. We evaluate the SelfAware dataset (Yin et al., 2023) using LLaMA3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, which consists of 2337 answerable and 1032 unanswerable questions from five diverse categories. Specifically, given a subjective question in SelfAware, we ask an LLM to evaluate the answerability of this question on a scale of 1 to 5. A score of 1 indicates that the question is completely unanswerable based on the available knowledge, while a score of 5 indicates a highly accurate answer is possible. Following the original paper, we normalize the certainty score to a range of 0-1 and use a threshold of 0.75 as the boundary to classify LLM as either “know” ($\geq$ 0.75) or “don’t know” (<< 0.75).
Then, we can quantitatively calculate the F1 score to measure the model’s level of self-knowledge.Table 8 compares lager with baseline methods, including the SimCSE baseline used in the original paper. The results show that lager (w. tuning) significantly enhances the LLM’s self-knowledge, achieving a +19.7% and +1.9% increase in F1 score compared to SimCSE, and VScore on LLaMA-3.1-8B-Instruct, respectively, and +11.1%, and +4.8% increases in F1 score compared to SimCSE and VScore on Mistral-7B-Instruct-v0.3, respectively. By providing a more reliable measure of LLM’s self-knowledge, lager enhances the LLM’s ability to understand their limitations, shows performance comparable to the E-score.
Appendix E Experiment Details
E.1 Details About KDE for Score Distribution Visualization
We adopt kernel density estimation (KDE) to visualize the distribution of model-generated scores across evaluation samples. KDE provides a smooth, non-parametric estimate of the underlying score distribution and is preferable to histograms for highlighting fine-grained density differences.
Given a set of samples $\{x_{1},x_{2},\dots,x_{n}\}$, the estimated density at point $x$ is computed as:
$\hat{f}_{h}(x)=\frac{1}{nh}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right),$ (14)
where $K(\cdot)$ is the kernel function and $h$ is the bandwidth. We use the Gaussian kernel:
$K(u)=\frac{1}{\sqrt{2\pi}}e^{-u^{2}/2}.$ (15)
All KDE plots are normalized to unit area for comparability. We employ kernel density estimation for visualization, with the bandwidth parameter selected using Silverman’s rule or cross-validation when appropriate. KDE is used in Figure 4 to compare the score output distributions of lager and baseline models.
E.2 Details About the Benchmarks
FLASK
We utilize the complete test prompt set from FLASK. (Ye et al., 2024a), a fine-grained evaluation dataset that includes various conventional NLP datasets and instruction datasets. This dataset contains data on 2,001 entries, each consisting of a human feedback score and an evaluation score from GPT-4. We utilize 12 scoring rubrics: Conciseness, Metacognition, Insightfulness, Readability, Commonsense Understanding, Logical Robustness, Factuality, Comprehension, Completeness, Logical Efficiency, Logical Correctness, and Harmlessness.
HelpSteer
HelpSteer (Wang et al., 2023) is an open-source Helpfulness Dataset. For each response in the dataset, it is evaluated based on five criteria: helpfulness, correctness, coherence, complexity, and verbosity. These evaluation scores are annotated by humans, ensuring that each instruction in the dataset is paired with responses of varying quality. We process these data, combining the prompt, response, and corresponding criteria into individual data samples. A total of 8.95k data points are generated, with the first 2k used for our evaluation.
BiGGen Bench
BiGGen Bench (Kim et al., 2024b) is a comprehensive evaluation benchmark designed to assess the capabilities of large language models (LLMs) across a wide range of tasks. This benchmark focuses on free-form text generation and employs fine-grained, instance-specific evaluation criteria. BiGGen Bench evaluates nine distinct generation capabilities (e.g., instruction following, reasoning, tool usage, etc.) across 77 tasks, providing model outputs and scores for 103 different language models. We utilize the human evaluation test set.
E.3 Details About the Baselines
Due to the limitations of the GPTScore method, we do not report its results under the reasoning condition. Since TIGERScore-7B and Prometheus2-7B cannot directly output a single score, we do not report their results under the direct condition.
TIGERScore
TIGERScore(Jiang et al., 2023b) is a trained metric designed for explainable, reference-free evaluation of text generation tasks. Unlike other automatic evaluation methods that only provide scores, TIGERScore uses natural language instructions to guide error analysis, identifying specific mistakes in the generated text. It is based on the LLaMA-2 model and trained on a carefully curated dataset, MetricInstruct, which includes 42K quadruples covering 6 text generation tasks and 23 datasets. TIGERScore is capable of generating detailed error analysis, including the error location, error type, error explanation, and revision suggestions, along with a penalty score for each error.
GPTScore
Assuming the text to be evaluated is $\bm{h}=\{h_{1},h_{2},\cdots,h_{m}\}$, The core of the GPTscore method lies in leveraging LLM to evaluate the quality of text by computing the average log probability $\frac{1}{m}\sum_{t=1}^{m}\log p(h_{t}|\bm{h_{<t}},T(d,a,S);\theta)$ under a specific context $S$ and task instruction $T(d,a,S)$. The prompt template $T(\cdot)$ is composed of a task description and aspect definition.
Prometheus-2-7B
The process of constructing the Prometheus-2 model (Kim et al., 2024b) involves the following steps: First, a fine-grained pairwise ranking dataset, PreferenceCollection, containing 1000 custom evaluation criteria, is created to support evaluation based on user-defined standards. Next, Mistral-7B is selected as the base model, and it is trained separately on the direct evaluation dataset FEEDBACK COLLECTION and the pairwise ranking dataset PreferenceCollection. Finally, by merging the weights of the models trained on these two formats, the final Prometheus-2 model is obtained.
E.4 Details of Evaluation Metrics
Pearson
Pearson correlation coefficient indicates the strength and direction of the relationship between two variables, ranging from -1 to 1. For the two sets of data x,y, we can calculate their Pearson correlation coefficients $\rho_{XY}$ as follows:
$\rho_{XY}=\frac{Cov(X,Y)}{\sigma_{X}\sigma_{Y}}=r_{p}$ (16)
Spearman
Spearman rank correlation coefficient is a non-parametric measure that evaluates the statistical dependence between the ranks of two variables. It determines how well the relationship between these variables can be described using a monotonic function. We can calculate Spearman $r_{s}$ according to the following formula:
$r_{s}=1-\frac{6\sum_{i=1}^{n}d^{2}}{n(n^{2}-1)}$ (17)
where $d_{i}$ represents the rank difference between the two variables for each observation (pairwise difference), and $n$ denotes the number of fraction pairs.
Since the Spearman correlation coefficient effectively measures nonlinear monotonic relationships, is robust to outliers, and does not require data to follow a normal distribution, it is particularly suitable for handling ordinal data. Therefore, this work primarily presents results based on the Spearman correlation coefficient.
E.5 Details of layer-wise weight training settings
We randomly select 1,000 samples from the HelpSteer dataset as a held-out validation set, ensuring that it does not overlap with the test set, to tune the layer-wise weights of lager. The training is performed using the Adam optimizer with an initial learning rate of 0.01, and a batch size of 4. A random seed of 42 is set to ensure the reproducibility of the experiment. We also apply the ReduceLROnPlateau learning rate scheduler with a decay factor of 0.5, a patience value of 1, and a minimum learning rate specified by min_lr. Since some models in the Qwen2.5 family may not converge with just one epoch, we set the number of training epochs to 2 for all models in the Qwen2.5 series. For other backbone models, we set the number of training epochs to 1. All experiments are implemented using PyTorch.
Appendix F Details of SFT
For lager, E-Score and VScore, we establish seven dimensions (including answer accuracy, logical consistency, relevance, fluency and clarity, length appropriateness, diversity, and instruction difficulty) and utilize the average score of each dimension, evaluated using the LLaMA3.1-8B-Instruct model, for data selection. The specific definitions and scoring standards for these seven dimensions can be found in Table 12. SuperFiltering is an explicitly designed instruction data filtering method. We strictly follow the original paper’s methodology, except for using LLaMA3.1-8B-Instruct as the backbone model for fair comparison. During the SFT training of the Llama-3-8B model, we use 4 L40 GPUs, and the total batch size during the training phase is 64. The specific training parameters and settings are detailed in Table  9.
Table 9: Parameters and settings used for instruction fine-tuning of the Llama-3-8B model.
Training Parameters Setting
stage sft
finetuning_type full
template alpaca
flash_attn fa2
cutoff_len 2048
learning_rate 2e-5
num_train_epochs 3
per_device_train_batch_size 4
gradient_accumulation_steps 4
lr_scheduler_type cosine
warmup_ratio 0.03
packing FALSE
bf16 TRUE
tf32 TRUE
optim adamw_torch
include_num_input_tokens_seen TRUE
seed 42
Appendix G Prompt Template
In this section, we present all the prompt templates used in this paper, with the blue text indicating the parts where corresponding content should be inserted. For the direct evaluation scenario, Figure 9 is the prompt template for direct evaluation in a point-wise scenario that outputs only a single score. Figure 10 is the prompt template for reasoning evaluation in a point-wise scenario, including feedback(reasoning).
Figure 8: The prompt template for Emotional Understanding experiment.
Figure 9: The prompt template for direct evaluation in a point-wise scenario that outputs only a single score.
Figure 10: The prompt template for reasoning evaluation in a point-wise scenario, including feedback(reasoning).
Figure 11: The prompt template used for data filtering in SFT task.
Figure 12: The prompt template used in the SelfAware dataset in section 7.2.
{longtblr}
[ caption = Definitions of different dimensions and their specific scoring standards, with each dimension scored on a scale ranging from 1 to 9, each score corresponds to a specific standard., ] width = colspec = Q[150]Q[812], row4 = t, column1 = c, cell21 = r=9, cell121 = r=9, cell221 = r=9, cell321 = r=9, cell421 = r=9, cell521 = r=9, cell621 = r=9, vlines, hline1,71 = -0.12em, hline2,11-12,21-22,31-32,41-42,51-52,61-62 = -, hline3-10,13-20,23-30,33-40,43-50,53-60,63-70 = 2, Dimension-1 & Definition  Answer Accuracy: Evaluate whether the response accurately addresses the instruction and completely fulfills the task.
Scoring Standards 1: Completely incorrect, irrelevant to the instruction.
2: Partially correct, major omissions or errors in fulfilling the instruction.
3: Contains significant errors, unable to fully address the core task.
4: Partially correct, missing key details or addressing the wrong aspect of the instruction.
5: Mostly accurate, but contains some errors or omissions.
6: Mostly correct, though missing small details or has minor inaccuracies.
7: Largely accurate and complete, but may lack small details or have minimal errors.
8: Fully accurate, completely addresses the instruction, minimal flaws.
9: Perfectly accurate, fully aligns with the instruction, no omissions.
Dimension-2 & Definition Logical Consistency: Assess whether the response maintains logical consistency, following the reasoning of the instruction without contradictions.
Scoring Standards 1: Completely incoherent, self-contradictory, does not follow the instruction.
2: Major logical errors or contradictions, does not follow the instruction.
3: Logic is unclear, unable to fully support the instruction’s requirements.
4: Contains some logical inconsistencies, but overall understandable, with minor deviation from the instruction.
5: Mostly consistent, but has slight logical flaws or areas that could be clearer.
6: Largely logical and clear, but small inconsistencies or areas for improvement exist.
7: Logical and coherent, only minimal inconsistencies remain.
8: Completely logical, reasoning is sound and aligns perfectly with the instruction.
9: Perfectly logical, tightly reasoned, flawless adherence to the instruction’s requirements.
Dimension-3 & Definition Relevance: Evaluate whether the response is relevant to the instruction, directly addressing the key aspects of the task.
Scoring Standards 1: Completely irrelevant, deviates from the instruction’s theme.
2: Minimal relevance, strays far from the core task.
3: Poor relevance, does not adequately address the main parts of the instruction.
4: Somewhat relevant, but deviates from key details or task elements of the instruction.
5: Mostly relevant, but lacking some precision or key aspects of the instruction.
6: Mostly relevant, covers the main aspects of the instruction, but can be improved for accuracy or detail.
7: Directly relevant, clearly and adequately responds to the instruction.
8: Highly relevant, fully addresses the core aspects of the instruction with precision.
9: Perfectly relevant, fully and comprehensively addresses all aspects of the instruction.
Dimension-4 & Definition Fluency Clarity: Evaluate the fluency and clarity of the response, ensuring it follows the expression requirements of the instruction and is easy to understand.
Scoring Standards 1: Completely unclear, frequent grammar errors, hard to understand.
2: Very unclear, poor structure, difficult to follow.
3: Unclear, sentence structure issues, parts are hard to understand.
4: Somewhat unclear, lacks fluency, but overall understandable.
5: Clear, but could be improved in flow or conciseness.
6: Clear and fluent, but slight improvements in clarity or conciseness could be made.
7: Fluent, easy to understand, clear expression.
8: Very clear and fluent, perfect adherence to the instruction’s expression requirements.
9: Perfectly fluent, natural, and clear, fully aligned with the instruction’s expectations.
Dimension-5 & Definition Length Appropriateness: Evaluate whether the response length is appropriate, providing sufficient information without being too brief or too long according to the instruction.
Scoring Standards 1: Extremely short or overly long, irrelevant to the instruction’s expected length.
2: Too short or too long, fails to meet the instruction’s requirements.
3: Too brief, lacks key details, or too lengthy with unnecessary information.
4: Slightly short or slightly long, lacks important details or includes redundancy.
5: Reasonable length, but could be more concise or include more details as per the instruction.
6: Length is appropriate, but could be refined by removing unnecessary parts or adding small details.
7: Length is well-suited, adequately covers the instruction’s main points without redundancy.
8: Perfect length, concise yet comprehensive, fully meets the instruction’s requirements.
9: Optimal length, precisely conveys all required details, no redundancy, aligns perfectly with the instruction.
Dimension-6 & Definition Diversity: Evaluate whether the response shows diversity in language, structure, and viewpoints, and avoids repetition or overly uniform expressions as required by the instruction.
Scoring Standards 1: Extremely repetitive, no variation, fails to meet the instruction’s diversity requirements.
2: Lacks diversity, repetitive or formulaic expression, no innovation.
3: Some diversity, but much of the content is repetitive or uniform.
4: Some variation, but significant repetition or lack of diverse viewpoints.
5: Some diversity, but certain sections are somewhat repetitive or conventional.
6: Largely diverse, offers different perspectives or expressions, but with minor repetition.
7: Strong diversity, rich in varied language and structure, aligns with the instruction’s expectations.
8: Very diverse, with significant creativity and variety in language and perspective.
9: Extremely creative, highly diverse, fully meets the instruction’s innovation and diversity requirements.
Dimension-7 & Definition Instruction Difficulty: Assess the complexity of the instruction, considering whether it requires deep reasoning, multiple steps, or specialized knowledge, and how well the response reflects the difficulty of the task.
Scoring Standards 1: Very simple instruction, requires only basic information with no reasoning.
2: Simple task, requires basic common knowledge or simple answers.
3: Slightly more complex, requiring some background knowledge or understanding of specific content.
4: Moderately complex, involving some reasoning or moderately difficult tasks.
5: Instruction requires multiple steps or specialized knowledge across domains.
6: Complex instruction requiring advanced reasoning or a wide range of knowledge.
7: Highly complex, requires deep reasoning or tasks that span multiple domains.
8: Very complex, involves multi-layered reasoning or highly specialized knowledge.
9: Extremely complex, requiring professional-level knowledge or intricate reasoning to fulfill.

### v3

@@
 Abstract
-The growing scale of evaluation tasks has led to the widespread adoption of automated evaluation using large language models, a paradigm known as "LLM-as-a-judge." However, improving its alignment with human preferences without complex prompts or fine-tuning remains challenging.
-In this work, motivated by preliminary findings that middle-to-upper layers encode semantically and task-relevant representations that are often more aligned with human judgments than the final layer, we propose lager, a lightweight and efficient framework for enhancing LLM-as-a-Judge alignment with human scoring, via internal representations.
-lager produces fine-grained judgment scores by aggregating cross-layer score-token logits and computing the expected score from a softmax-based distribution, with the LLM backbone kept frozen.
+The growing scale of evaluation tasks has led to the widespread adoption of automated evaluation using LLMs, a paradigm known as "LLM-as-a-judge".
+However, improving its alignment with human preferences without complex prompts or fine-tuning remains challenging.
+Previous studies mainly optimize based on shallow outputs, overlooking rich cross-layer representations.
+In this work, motivated by preliminary findings that middle-to-upper layers encode semantically and task-relevant representations that are often more aligned with human judgments than the final layer, we propose lager, a post-hoc, plug-and-play framework for improving the alignment of LLM-as-a-Judge point-wise evaluations with human scores, by leveraging internal representations.
+lager produces fine-grained judgment scores by aggregating cross-layer score-token logits and computing the expected score from a softmax-based distribution, while keeping the LLM backbone frozen and ensuring no impact on the inference process.
 lager fully leverages the complementary information across different layers, overcoming the limitations of relying solely on the final layer.
@@
 Without reasoning steps, lager matches or outperforms reasoning-based methods.
-Experiments on downstream applications, such as data selection and emotional understanding, further show the effectiveness of our method.
+Experiments on downstream applications, such as data selection and emotional understanding, further show the generalization of lager.
 1 Introduction
@@
 lager leverages the semantic diversity across different layers, allowing the final score to integrate both low-level lexical cues and high-level reasoning signals, thereby providing a more comprehensive reflection of the model’s understanding of the scoring task.
-Unlike the finetuning-based method, the parameters of the general LLM backbone are always kept frozen.
+lager is plug-and-play, leaving model parameters and next-token predictions unchanged.
 lager outperforms baselines on alignment benchmarks such as Flask, HelpSteer, and BIGGen, with improvements of up to 7.5% in Spearman correlation.
@@
 Experiments on more downstream applications show the effectiveness of the lager method.
-[^1] Our code will be publicly available at https://github.com/sustech-nlp/LAGER.
+[^1] Our code is publicly available at https://github.com/sustech-nlp/LAGER.
 Figure 1: (a): Vanilla scores may overlook meaningful distinctions by relying solely on the most probable score token.
@@
 We denote the full input as $X=\{x_{d},x_{i},x_{r},x_{c}\}$, and the set of candidate scores as $\mathbb{S}\subset\mathcal{V}$, where $\mathcal{V}$ is the model’s vocabulary.
-Let $\mathbf{h}^{(L)}_{n}$ denote the final-layer hidden state at the position where the score token is to be generated (e.g., after ‘Score:’).
+Let $\mathbf{h}^{(L)}_{n}$ denote the final-layer hidden state at the position where the score token is to be generated (e.g., after ’Score:’).
 The model produces output logits:
@@
 by taking the expectation over the score distribution, we can obtain the final fine-grained judge score:
-$s^{*}=\mathbb{E}_{s\sim P(s)}[s]=\sum_{s\in\mathbb{S}}s\times P(s)$ (12)
+$s^{*}=\mathbb{E}_{s\sim P(s)}[s]=\sum_{s\in\mathbb{S}}s\times P(s).$ (12)
 Normalizing before aggregation removes the relative scale information of logits and overemphasizes less important layers. We compare both settings in the ablation study (Section 4.3).
@@
 One is to apply average aggregation $w_{l}=1/(L+1)$ (denoted as lager (w.o.
-tuning) in Table 1), the other is to tune the lightweight $L+1$ parameters with a small-scale validation set (denoted as lager (w.
+tuning) in Table 1), the other is to tune the lightweight $L+1$ parameters ($L$ is the number of transformer layers) with a small-scale validation set (denoted as lager (w.
 tuning)).
@@
 Please refer to Appendix E.5 for specific training settings and details.
-It is important to note that the weights are tuned only once for each backbone and subsequently applied across all benchmarks.
+It is important to note that the weights are tuned only once for each backbone and subsequently applied across all benchmarks and downstream tasks.
 As the LLM remains frozen, no additional judge model is needed for scenarios like self-improvement.
+Our method is implemented without altering the model’s next-token prediction process or logits, thereby offering plug-and-play capability.
 In contrast, finetuning-based methods require domain-specific judge models and suffer from limited generalization due to scarce labeled data.
 3.4 Discussion: Applicability under Restricted Access Settings
-While our proposed framework relies on accessing the intermediate layers of the LLM, it can naturally adapt to more restricted scenarios.
-For instance, when only final-layer logits are available (e.g., API-based models), we can extract the logits corresponding to score tokens, apply softmax, and compute the expected score.
-The method proposed in prior work (Liu et al., 2023) generates a continuous score estimate by avoiding argmax on discrete tokens.
+While lager relies on accessing the intermediate layers of the LLM, it can naturally adapt to more restricted scenarios.
+For instance, when only final-layer logits are available (e.g., API-based models), we can extract the logits corresponding to score tokens, apply softmax, and compute the expected score (this method is proposed in G-Eval (Liu et al., 2023)).
 We refer to this setup as expectation score (E-Score).
@@
 We treat both scores as comparable baselines in our experiments.
+Nonetheless, powerful open-source models still remain the mainstream for large-scale evaluation due to their lower cost and performance comparable to closed-source models (Lambert et al., 2024; Gu et al., 2024; Malik et al., 2025), making lager broadly applicable and easily extensible to future models.
 4 Experiments
 4.1 Experiments Setup
-Benchmarks
+Benchmarks.
 Due to potential preference leakage (Li et al., 2025) between models, we evaluate our method on human-annotated datasets for reliable results.
-We chose three point-wise benchmarks to ensure a comprehensive evaluation: Flask (Ye et al., 2024a), HelpSteer (Wang et al., 2023), and BiGGen Bench [^3] https://huggingface.co/datasets/prometheus-eval/BiGGen-Bench-Results  (Kim et al., 2024b).
+We chose three diverse point-wise benchmarks to ensure a comprehensive evaluation: Flask (Ye et al., 2024a), HelpSteer (Wang et al., 2023), and BiGGen Bench [^3] https://huggingface.co/datasets/prometheus-eval/BiGGen-Bench-Results  (Kim et al., 2024b).
 In these benchmarks, human-annotated scores range from 1 to 5.
 Please refer to Appendix E.2 for the details.
-Models
+Models.
 We experiment with multiple backbone models of varying sizes and from different families, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), InternLM3-8B-Instruct (Cai et al., 2024), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), Qwen-2.5-14B-Instruct (Qwen et al., 2025), Mistral-Small-24B-Instruct and LLAMA-3.3-70B-Instruct.
-Baselines
+Baselines.
 We compare lager with three baselines: GPTScore, vanilla score (VScore) and expectation score (E-Score).
@@
 Please refer to Appendix E.3 for specific details about the baselines.
-Evaluation Metrics
+Evaluation Metrics.
 Following previous works (Gu et al., 2024; Bai et al., 2023; Liu et al., 2025), we use the commonly adopted Pearson and Spearman correlations to measure the consistency between LLM-predicted scores and human annotations.
 See Appendix E.4 for further details.
-Evaluation details
-We adopt greedy decoding for its stability and consistency in evaluation (Lin & Chen, 2023).
-To comprehensively assess the effectiveness of our method, we conduct evaluations under two different conditions: 1) Direct: Directly providing a score during evaluation without reasoning.
-2) Reasoning: Previous work (Liu et al., 2023; Wang et al., 2024a; Hu et al., 2024a) shows that reasoning improves judgment quality, with reasoning-first setups performing best(Chen et al., 2024; Murugadoss et al., 2024; Gu et al., 2024).
-Thus, we adopt a reasoning-first approach in this study.
-Please refer to Appendix G for the prompt templates.
-Unless otherwise specified, we highlight the best results in each table using bold, and denote the second-best with underlining.
-Table 1: Spearman Correlation Results (Part 1 of 2): lager vs.
+Table 1: Spearman Correlation Results: lager vs.
 Baselines on Flask, HelpSteer, and BiGGen Bench.
-See Table 5 for the remaining results.
 These results are statistically significant, with p-values less than 1e-5.
@@
 Open-source Models
+Mistral-7B-Instruct-v0.3
+GPTScore 0.258 - 0.209 - 0.183 - 0.217
+Vscore 0.266 0.269 0.267 0.364 0.138 0.280 0.264
+E-Score 0.239 0.279 0.296 0.380 0.185 0.283 0.277
+lager (w.o tuning) 0.338 0.295 0.401 0.377 0.353 0.329 0.349
+lager (w. tuning) 0.347 0.298 0.403 0.376 0.357 0.333 0.352
+LLaMA3.1-8B-Instruct
+GPTScore 0.061 - -0.022 - -0.162 - -0.041
+Vscore 0.334 0.429 0.374 0.518 0.273 0.390 0.386
+E-Score 0.386 0.446 0.464 0.525 0.352 0.403 0.429
+lager (w.o tuning) 0.472 0.456 0.520 0.524 0.475 0.443 0.482
+lager (w. tuning) 0.477 0.460 0.515 0.524 0.482 0.444 0.484
 InternLM3-8B-Instruct
@@
 lager (w.o tuning) 0.449 0.468 0.426 0.429 0.374 0.469 0.436
-lager (w. tuning) 0.545 0.489 0.515 0.474 0.507 0.49 0.501
+lager (w. tuning) 0.545 0.489 0.515 0.474 0.507 0.490 0.501
 Qwen-2.5-14B-Instruct
@@
 lager (w. tuning) 0.596 0.542 0.449 0.487 0.598 0.566 0.540
+LLAMA-3.3-70B-Instruct
+GPTScore 0.042 - 0.014 - -0.193 - -0.046
+Vscore 0.518 0.567 0.435 0.494 0.559 0.539 0.519
+E-Score 0.464 0.540 0.444 0.488 0.530 0.506 0.495
+lager (w.o tuning) 0.610 0.598 0.506 0.520 0.597 0.585 0.569
+lager (w. tuning) 0.611 0.598 0.504 0.519 0.602 0.584 0.570
+Evaluation Details.
+We adopt greedy decoding for its stability and consistency in evaluation (Lin & Chen, 2023).
+To comprehensively assess the effectiveness of our method, we conduct evaluations under two different conditions: 1) Direct: Directly providing a score during evaluation without reasoning.
+2) Reasoning: Previous work (Liu et al., 2023; Wang et al., 2024a; Hu et al., 2024a) shows that reasoning improves judgment quality, with reasoning-first setups performing best(Chen et al., 2024; Murugadoss et al., 2024; Gu et al., 2024).
+Thus, we adopt a reasoning-first approach in this study.
+The prompt template used in our work follows the standard configuration widely adopted in prior studies (Kim et al., 2024b; Liu et al., 2023; Xu et al., 2023), please refer to Appendix G for the prompt templates.
+Unless otherwise specified, we highlight the best results in each table using bold, and denote the second-best with underlining.
 4.2 Main Results
-Tables 1 and 5 report Spearman results of lager compared to baselines across three benchmarks.
+Tables 1 report Spearman results of lager compared to baselines across three benchmarks.
 Section 5.3 further analyzes the performance variation of lager and baseline methods across different model scales, under the setting where models from the Qwen2.5 family are used as frozen backbones, and compares their results on the same set of benchmarks.
-The Effectiveness and Scalability of LAGER
+Specifically, lager (w.
+tuning) achieves near-best performance on these benchmarks and demonstrates strong stability, so it serves as the main approach in our work.
+The Effectiveness and Scalability of LAGER.
 lager with tuning consistently outperforms all non-training baselines across diverse LLM backbones on all three benchmarks.
@@
 Due to its focus on error detection, the fine-tuned TIGERscore-7B shows relatively limited performance.
-Compared to the fine-tuned Prometheus2-7B, lager also enables InternLM3-8B-Instruct and LLaMA3.1-8B-Instruct(results shown in Table 5), two models of comparable size, to surpass it even though they initially have weaker performance on Vscore.
-The Robustness of LAGER with Tuning
+Compared to the fine-tuned Prometheus2-7B, lager also enables InternLM3-8B-Instruct and LLaMA3.1-8B-Instruct, two models of comparable size, to surpass it even though they initially have weaker performance on Vscore.
+Even without tuning, lager outperforms E-Score in 5 out of 6 backbone models, with an average improvement of 4.5%.
+In the few cases where it underperforms, the performance gap is within 3%.
+Moreover, it consistently outperforms V-Score.
+The Robustness of LAGER.
 The average E-Score performance of models like Qwen-2.5-14B-Instruct, LLAMA-3.3-70B-Instruct, and Mistral-7B-Instruct-v0.3 is sometimes lower than VScore.
-Specifically, the average E-score performance of LLAMA-3.3-70B-instruct (results shown in Table 5) across all benchmarks is 2.4 points lower than Vscore.
+Specifically, the average E-score performance of LLAMA-3.3-70B-instruct across all benchmarks is 2.4 points lower than Vscore.
 In contrast, lager with tuning achieves almost the best performance across all settings and benchmarks, consistently outperforms Vscore.
@@
 More details are discussed in Section 5.
-Reasoning Is Not Always Better
+Therefore, we primarily recommend the lager with tuning approach and conduct subsequent analyses based on it.
+The performance gains of LAGER stem not from large-scale fine-tuning, but from a more efficient and structured utilization of the model’s internal representations.
+Reasoning Is Not Always Better.
 The richness of information aggregation increases gradually from Vscore, E-Score to lager (with tuning).
@@
 We speculate that this phenomenon may be due to reasoning potentially causing the model to become overconfident (Yang et al., 2024; Wagner et al., 2024), which in turn leads to inaccurate evaluation scores.
-4.3 Ablation Study
+This is also consistent with the findings of previous work (Ge et al., 2025; Guo et al., 2024; Li et al., 2024d): As the reasoning steps progress, the model’s focus on the original input and the text to be evaluated gradually diminishes, shifting more toward its self-generated reasoning trajectory.
 Table 2: Ablation Study Results on FLASK and HelpSteer Datasets Evaluated by Spearman Corr.
@@
 […]
-To better understand each design choice of lager, we conduct a comprehensive ablation study on the direct setup using Mistral-7B-Instruct-v0.3 and InternLM3-8B-Instruct models.
+4.3 Ablation Study
+To better understand each design choice of lager, we conduct a comprehensive ablation study on the direct setup using Mistral-7B-Instruct-v0.3 and InternLM3-8B-Instruct.
 Table 2 shows the results of the Spearman correlation.
-The expectation (Exp.) score and the maximum (Max.) score are two alternative designs.
+The expectation (Exp.) score and the maximum (Max.) score are alternative designs.
 Given the score distribution of a certain layer, we use Eqn.12 to calculate the expected score or use the score with the highest probability.
@@
 Multi-layer integration is generally effective, especially for Mistral; however, in untuned settings, InternalLM occasionally performs better without aggregation, indicating model-specific differences.
+5 Analyses
+5.1 Understanding Internal States
 Figure 3: Analysis of Llama3.1-8B-Instruct and Mistral-7B-Instruct-V0.3 using HelpSteer and Flask.
@@
 (c) depicts the average cosine similarity heatmap of hidden states across layers.
-5 Analyses
-5.1 Understanding Internal States
 Having observed that internal states contribute significantly to the success of lager, we now aim to understand how internal states in each layer contribute to the judgment score.
@@
 lager (w. tuning) 0.087 0.060
-We also quantitatively compare lager, E-Score and VScore by evaluating the Kullback-Leibler (KL) divergence and the mean squared error (MSE).
+We also quantitatively compare lager, E-Score, and VScore by evaluating the Kullback-Leibler (KL) divergence and the mean squared error (MSE).
 The KL divergence of a candidate score $S$ is computed as $\mathrm{D}_{\mathrm{KL}}(\tilde{p}_{hs}||\tilde{p}_{s})$ where $\tilde{p}_{hs}$ and $\tilde{p}_{s}$ represent the human score distribution and the candidate score distribution, respectively.
-The results, presented in Table 3, demonstrate that lager more accurately approximates the distribution of human-annotated scores.
+The results in Table 3, demonstrate that lager more accurately approximates the distribution of human-annotated scores.
 5.3 From Tiny to Huge: LAGER Brings Improvements Across Model Scales
@@
 Notably, on the BIGGen dataset, lager shows a significant improvement on the 0.5B model scale, elevating the Spearman correlation from a negative value (-0.4 for the Vanilla Score) to a positive 0.1, marking a substantial gain compared to the baseline.
-These results demonstrate its effectiveness across different model scales and datasets.
+These results demonstrate its effectiveness across different model scales and datasets, and the benefits brought by lager are scale-transferable and synergistically amplified as model capacity grows.
 Figure 5: Comparison of Spearman Results for Qwen2.5 Models Across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) (Direct Condition).
@@
 In this experiment, we use lager as a scoring metric to select a 10% subset from the alpaca-cleaned-52k dataset [^4] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023).
-We compare our subset, Highest lager, with four baselines on finetuning the LLaMA3-8B-base model: Longest instructions, Longest responses (Shen, 2024), Highest VScore, and SuperFiltering (Li et al., 2024d).
+We compare our subset, Highest lager, with four baselines on finetuning the LLaMA3-8B-base model: Longest instructions, Longest responses (Shen, 2024), Highest VScore, and SuperFiltering (Li et al., 2024e).
 SuperFiltering is an explicitly designed instruction data filtering method.
@@
 lager is quite general and can be applied to various use cases, including but not limited to instruction data selection, recognizing its knowledge boundaries and emotional understanding.
+Acknowledgements
+This project was supported by National Natural Science Foundation of China (No.
+62306132), Guangdong Basic and Applied Basic Research Foundation (No.
+2025A1515011564), Natural Science Foundation of Shanghai (No.
+25ZR1402136).
+We thank the anonymous reviewers for their insightful feedback on this work.
 Appendix A Impact Statement
@@
 G-Eval (Liu et al., 2023) evaluates NLG output quality using Chain-of-Thought (CoT) and Form-Filling paradigms.
-Portia (Li et al., 2024e) proposes to evaluate the responses with step-by-step reasoning guidelines.
+Portia (Li et al., 2024f) proposes to evaluate the responses with step-by-step reasoning guidelines.
 Active-critic (Xu et al., 2024) first generates evaluation criteria from data and iteratively refines them.
@@
 tuning) under the reasoning condition.
-We present the Spearman correlation coefficients of lager and other baselines for LLaMA3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Mistral-Small-24B-Instruct on Flask, HelpSteer, and BiGGen Bench, as shown in Table 5.
 In Table 5, we present the Pearson correlation coefficients for all models, including lager and other baselines, on the Flask, HelpSteer, and BiGGen Bench.
@@
 Figure 7: Comparison of Spearman Correlation Coefficients for Qwen2.5 Models Across FLASK, HelpSteer and BIGGen Bench Using Vanilla Score, E-Score, and lager (w. tuning) (Reasoning Condition)
-Table 5: Spearman Correlation Results (Part 2 of 2): Continuation of Table 1.
-These results are statistically significant, with p-values less than 1e-5.
-See Table 6 for the full Pearson correlation results.
-Model
-Flask
-Direct
-Flask
-Reasoning
-HelpSteer
-direct
-HelpSteer
-Reasoning
-BIGGen Bench
-Direct
-BIGGen Bench
-Reasoning
-Average
-Fine-tuned Models
-TIGERscore-7B - 0.175 - 0.118 - 0.171 0.155
-Prometheus2-7B - 0.413 - 0.514 - 0.389 0.448
-Close-source Model via API
-GPT-4o-mini
-Vscore 0.526 0.535 0.482 0.535 0.534 0.509 0.520
-E-Score 0.579 0.561 0.500 0.541 0.573 0.530 0.547
-Open-source Models
-LLaMA3.1-8B-Instruct
-GPTScore 0.061 - -0.022 - -0.162 - -0.041
-Vscore 0.334 0.429 0.374 0.518 0.273 0.390 0.386
-E-Score 0.386 0.446 0.464 0.525 0.352 0.403 0.429
-lager (w.o tuning) 0.472 0.456 0.520 0.524 0.475 0.443 0.482
-lager (w. tuning) 0.477 0.460 0.515 0.524 0.482 0.444 0.484
-Mistral-7B-Instruct-v0.3
-GPTScore 0.258 - 0.209 - 0.183 - 0.217
-Vscore 0.266 0.269 0.267 0.364 0.138 0.280 0.264
-E-Score 0.239 0.279 0.296 0.380 0.185 0.283 0.277
-lager (w.o tuning) 0.338 0.295 0.401 0.377 0.353 0.329 0.349
-lager (w. tuning) 0.347 0.298 0.403 0.376 0.357 0.333 0.352
-LLAMA-3.3-70B-Instruct
-GPTScore 0.042 - 0.014 - -0.193 - -0.046
-Vscore 0.518 0.567 0.435 0.494 0.559 0.539 0.519
-E-Score 0.464 0.540 0.444 0.488 0.530 0.506 0.495
-lager (w.o tuning) 0.610 0.598 0.506 0.520 0.597 0.585 0.569
-lager (w. tuning) 0.611 0.598 0.504 0.519 0.602 0.584 0.570
 Table 5: Pearson correlation coefficients of lager and other baselines on Flask, HelpSteer, and BiGGen Benchmarks.
@@
 D.2 Enhancing LLM in Knowing When They Don’t Know
+Knowing when they don’t know, also known as self-knowledge (Yin et al., 2023), is crucial for enhancing the reliability and trustworthiness of LLMs.
+In this section, we explore whether lager can improve an LLM’s self-knowledge.
+We evaluate the SelfAware dataset (Yin et al., 2023) using LLaMA3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, which consists of 2337 answerable and 1032 unanswerable questions from five diverse categories.
+Specifically, given a subjective question in SelfAware, we ask an LLM to evaluate the answerability of this question on a scale of 1 to 5.
+A score of 1 indicates that the question is completely unanswerable based on the available knowledge, while a score of 5 indicates a highly accurate answer is possible.
+Following the original paper, we normalize the certainty score to a range of 0-1 and use a threshold of 0.75 as the boundary to classify LLM as either "know" ($\geq$ 0.75) or "don’t know" (<< 0.75).
 Table 7: Self-knowledge Comparison of lager and baselines. The evaluation metric is the F1-Score.
@@
 LLaMA-3.1-8B-Instruct 0.490 0.553 0.574 0.59 0.570
-Knowing when they don’t know, also known as self-knowledge (Yin et al., 2023), is crucial for enhancing the reliability and trustworthiness of LLMs.
-In this section, we explore whether lager can improve an LLM’s self-knowledge.
-We evaluate the SelfAware dataset (Yin et al., 2023) using LLaMA3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, which consists of 2337 answerable and 1032 unanswerable questions from five diverse categories.
-Specifically, given a subjective question in SelfAware, we ask an LLM to evaluate the answerability of this question on a scale of 1 to 5.
-A score of 1 indicates that the question is completely unanswerable based on the available knowledge, while a score of 5 indicates a highly accurate answer is possible.
-Following the original paper, we normalize the certainty score to a range of 0-1 and use a threshold of 0.75 as the boundary to classify LLM as either “know” ($\geq$ 0.75) or “don’t know” (<< 0.75).
 Then, we can quantitatively calculate the F1 score to measure the model’s level of self-knowledge.Table 7 compares lager with baseline methods, including the SimCSE baseline used in the original paper.
@@
 During the SFT training of the Llama-3-8B model, we use 4 L40 GPUs, and the total batch size during the training phase is 64.
-The specific training parameters and settings are detailed in Table  9.
+The specific training parameters and settings are detailed in Table  8.
 Table 8: Parameters and settings used for instruction fine-tuning of the Llama-3-8B model.